OpenStudy (anonymous):

simplify (√7+√5)^2 using Foil.......multiply the first terms what are the products

OpenStudy (watchmath):

:)

OpenStudy (anonymous):

smiles made me smile lol

OpenStudy (anonymous):

\[(\sqrt{7}+\sqrt{5})(\sqrt{7}+\sqrt{5}) = \sqrt{7}^2+2\sqrt{35}+\sqrt{5}^2\] \[12 + 2\sqrt{35}\]

OpenStudy (gw2011):

I hope this is what you are looking for: square root 7*square root 7=7 square root 7*square root 5= square root 35 square root 5*square root 7=square 35 square root 5*square root 5=5 Therefore, the answer is 12+2(square root 35)

OpenStudy (anonymous):

^ + 1 for writing by hand

OpenStudy (gw2011):

Correction: square 35 should be square root 35

OpenStudy (anonymous):

What's Foil?!